# Learn by Reasoning: Analogical Weight Generation for Few-Shot Class-Incremental Learning

### v1

@@ IV-E Ablation Study
To better understand the design and effectiveness of the proposed Brain-Inspired Analogical Generator (BiAG), we conduct extensive ablation studies on three FSCIL benchmark datasets. We focus on evaluating the contributions of the core components, the effect of different network depths, and training configurations. The overall results are illustrated in Fig. 4 and detailed in Table IX.
IV-E1 The Effectiveness of Each Component
@@
We evaluate the effectiveness of three key modules in BiAG: the Weight & Prototype Analogical Attention (WPAA), the Weight Self-Attention (WSA), and the Semantic Conversion Module (SMC). As illustrated in Fig. 4(a), the addition of each component leads to a consistent and progressive improvement in final session accuracy across all three benchmarks, confirming their complementary roles.
Among the three, WPAA contributes the most significant performance gain. It provides a mechanism to establish analogical associations between new class prototypes and previously learned semantic structures, enabling effective weight generation even under limited supervision. As shown in Fig. 4(a), adding WPAA improves the final session accuracy by 1.70% on CUB-200, 1.61% on miniImageNet, and 2.68% on CIFAR-100 over the base model. However, when used in isolation, the improvements diminish in later sessions, suggesting limited temporal stability. Adding WSA on top of WPAA introduces weight-level relational modeling, which helps the model maintain more consistent predictions as the incremental process progresses. The performance benefits accumulate further with the inclusion of SMC, which enhances the semantic compatibility between the feature space and the classifier space. When all three modules are incorporated, BiAG obtains the highest final session accuracy across all benchmarks, achieving cumulative improvements of 1.90% on CUB-200, 2.26% on miniImageNet, and 3.03% on CIFAR-100.
@@
The session-wise analysis in Fig. 4(b) reveals that WPAA significantly improves early-stage performance but suffers from growing degradation without additional modules. WSA alleviates this issue by reinforcing intra-weight consistency across sessions. SMC further stabilizes the learning process by aligning high-level semantics between features and weights, especially under growing distributional shifts. This trend is also supported by the results on CIFAR-100 in Table IX. The average accuracy increases from 62.21% (base) to 65.44% with WPAA, 66.70% with WSA, and 68.93% with all modules enabled—corresponding to an overall improvement of 6.72%. In the final session, the accuracy grows from 54.92% to 57.95%, confirming the effectiveness of our full analogical generation pipeline.
@@
These findings confirm that all three modules play essential and complementary roles. WPAA is fundamental for analogical generation, WSA stabilizes weight interaction over time, and SMC improves semantic transfer. Together, they lead to more robust and generalizable weight generation throughout the incremental learning process.
IV-E2 The Influence of the Number of Layers
We further investigate the effect of BiAG’s model depth by varying the number of stacked layers on the CUB-200 dataset. As shown in Fig. 4(c), increasing the number of layers leads to consistent improvements in final session accuracy. Specifically, the two-layer variant achieves 63.35%, while the four-layer and six-layer models reach 63.72% and 63.79%, respectively. These results suggest that deeper designs help BiAG capture richer analogical patterns and better align new class weights with prior knowledge. However, the performance gain from four to six layers is marginal, indicating diminishing returns as the depth increases. This trend suggests that a four-layer BiAG offers a favorable balance between model capacity and computational efficiency, capturing most of the essential cross-class relational structure without overcomplicating the architecture.
@@ IV-E2 The Influence of the Number of Layers
Figure 5: t-SNE visualization on CIFAR-100. (a) Base session: clear class clusters aligned with classifier weights. (b) Incremental session: BiAG-generated weights (red) better match new class features than prototype-based weights (black), preserving structure and improving generalization.
IV-F Visualization
@@
In Fig. 5, we visualize the feature distribution learned by our model using t-SNE [60] on the CIFAR-100 dataset. These visualizations offer an intuitive understanding of how BiAG organizes feature space across base and incremental sessions and how the generated weights help the model adapt to novel classes while retaining base knowledge.
Fig. 5(a) presents the feature representations after training on the base session. Each point corresponds to a sample, and different colors represent different base classes. The yellow triangles indicate the classifier weights learned for the base classes. As shown, samples of the same class are tightly clustered around their corresponding weights, forming clear and well-separated class boundaries. This indicates that the feature extractor successfully learns discriminative embeddings for base categories.
Fig. 5(b) shows the representations after an incremental session. In this case, black triangles denote the weights directly computed from new class prototypes, while red triangles denote the weights generated by BiAG. Notably, the generated weights (red) are more semantically aligned with the corresponding class clusters compared to the prototype-derived weights (black), which often deviate from the actual feature centers. This illustrates that BiAG can effectively leverage relational knowledge from old classes to generate more accurate and generalizable weights for novel classes.
Moreover, BiAG-generated weights not only improve the alignment of new class representations but also preserve the structure of the base class distribution. There is no significant shift or collapse in the base class clusters after incremental learning, demonstrating the model’s resistance to catastrophic forgetting. This validates the effectiveness of our analogical generation mechanism in maintaining inter-class separation and ensuring stable learning across sessions. The visualization highlights how BiAG enables balanced and robust incremental learning by bridging semantic gaps between base and novel classes in the classifier space.
@@ V Conclusion and Future Work
In this paper, we proposed the Brain-Inspired Analogical Generator (BiAG), a novel framework for Few-Shot Class-Incremental Learning (FSCIL). Drawing inspiration from human analogical reasoning, BiAG generates classification weights for novel classes by establishing semantic correspondences with previously learned knowledge—without requiring any parameter updates during incremental sessions. This design makes BiAG well-suited for data-scarce, dynamic learning scenarios. BiAG leverages old class weights, new prototypes, and a learnable query to produce semantically aligned weights via its three key components: WPAA, WSA, and SMC. These modules work together to improve generalization and reduce forgetting. Extensive experiments on CIFAR-100, miniImageNet, and CUB-200 show that BiAG consistently outperforms state-of-the-art methods across all key metrics, including average and final session accuracy.
In future work, we aim to explore the extension of this analogical generation paradigm to more complex continual learning scenarios. One direction is domain-incremental learning, where data from different distributions arrive sequentially, posing domain-shift challenges. Incorporating domain-aware analogical reasoning may enhance the model’s adaptability and robustness across heterogeneous environments. Additionally, integrating BiAG with more advanced backbone architectures or vision-language pretraining models (e.g., CLIP) may further enhance its semantic alignment capabilities, enabling few-shot incremental learning in open-vocabulary or multimodal settings. In summary, BiAG provides a new perspective on few-shot incremental learning through analogical reasoning. We believe this brain-inspired direction can inspire broader advancements in building flexible and cognitively motivated continual learning systems.

### v2

@@ IV-F Ablation Study
To better understand the design and effectiveness of the proposed Brain-Inspired Analogical Generator (BiAG), we conduct extensive ablation studies on three FSCIL benchmark datasets. We focus on evaluating the contributions of the core components, the effect of different network depths, and training configurations. The overall results are illustrated in Fig. 4 and detailed in Table X.
Figure 4: The results of ablation experiments. Ablation studies on the proposed BiAG. (a) Final session accuracy on three datasets after incrementally adding key components (WPAA, WSA, SCM); (b) Session-wise accuracy comparison on CUB-200, showing the contribution of each component throughout the incremental process; (c) Accuracy comparison of BiAG with different numbers of layers, demonstrating that a deeper design improves generalization to new classes.
IV-F1 The Effectiveness of Each Component
@@
We evaluate the effectiveness of three key modules in BiAG: the Weight & Prototype Analogical Attention (WPAA), the Weight Self-Attention (WSA), and the Semantic Conversion Module (SCM). As illustrated in Fig. 4(a), the addition of each component leads to a consistent and progressive improvement in final session accuracy across all three benchmarks, confirming their complementary roles.
[…]
@@
The session-wise analysis in Fig. 4(b) reveals that WPAA significantly improves early-stage performance but suffers from growing degradation without additional modules. WSA alleviates this issue by reinforcing intra-weight consistency across sessions. SCM further stabilizes the learning process by aligning high-level semantics between features and weights, especially under growing distributional shifts. This trend is also supported by the results on CIFAR-100 in Table X. The average accuracy increases from 62.44% to 65.11% with WPAA, 67.21% with WSA, and 68.93% with all modules enabled—corresponding to an overall improvement of 6.49%. In the final session, the accuracy grows from 54.92% to 57.95%, confirming the effectiveness of our full analogical generation pipeline.
@@
These findings confirm that all three modules play essential and complementary roles. WPAA is fundamental for analogical generation, WSA stabilizes weight interaction over time, and SCM improves semantic transfer. Together, they lead to more robust and generalizable weight generation throughout the incremental learning process.
@@
Figure 5: t-SNE visualization on CIFAR-100 comparing BiAG (Ours) and SVAC. (a) and (c) Base session: clear clusters aligned with classifier weights (yellow). (b) and (d) Incremental session: BiAG-generated weights (red) better align with new class features than prototype-based weights (black), while preserving the structure of base classes.
IV-F2 The Influence of the Number of Layers
We further investigate the effect of BiAG’s model depth by varying the number of stacked layers on the CUB-200 dataset. As shown in Fig. 4(c), increasing the number of layers leads to consistent improvements in final session accuracy. Specifically, the two-layer variant achieves 63.35%, while the four-layer and six-layer models reach 63.72% and 63.79%. These results suggest that deeper designs help BiAG capture richer analogical patterns and better align new class weights with prior knowledge. However, the performance gain from four to six layers is marginal, indicating diminishing returns as the depth increases. This trend suggests that a four-layer BiAG offers a favorable balance between model capacity and computational efficiency, capturing most of the essential cross-class relational structure without overcomplicating the architecture.
Figure 6: Visualization of analogical reasoning. We show the base class and the top-3 most similar classes based on the generated weights.
TABLE XII: Comparison of GPU Memory Usage, Runtime Efficiency, and Average Accuracy Across Different Training Stages on CUB-200
Method Base Training BiAG Training Incremental Training Total Avg
Memory Epochs Time Memory Epochs Time Memory Epochs Time Time Acc
(GB) (min) (GB) (min) (GB) /Each Session (min) (min) (%)
NC-FSCIL 1.45 80 56.60 1.45 105–150 3.97 60.57 67.28
SVAC 18.93 100 45.36 10.81 10 4.59 49.95 69.35
YourSelf - - - 9.51 100 212.5 212.5 69.85
Ours 3.44 200 23.70 1.28 200 10.80 1.33 1 2.10 36.60 70.71
IV-G Visualization
IV-G1 Feature Structure
@@
Figure 5 presents a t-SNE [60] visualization of feature distributions learned on CIFAR-100, comparing BiAG (Ours) and the baseline SVAC. These visualizations provide an intuitive understanding of how BiAG organizes the feature space across base and incremental sessions and how the generated weights adapt to novel classes while retaining base knowledge.
Figures 5(a) and (c) show the base session results. Each point corresponds to a sample, and different colors represent different base classes. Yellow triangles indicate the classifier weights for the base classes. Both methods form clear and well-separated clusters.
Figures 5(b) and (d) illustrate the incremental session results. Black triangles denote weights computed from novel class prototypes, while red triangles represent BiAG-generated weights. Compared to SVAC, BiAG-generated weights align more closely with the corresponding novel class clusters and preserve the base class structure without significant drift. This demonstrates that BiAG can leverage analogical reasoning from old classes to produce accurate and generalizable weights for new classes, which are more representative of the class and less affected by outliers, thus avoiding significant drift.
Overall, these visualizations show that BiAG maintains inter-class separation, reduces the risk of catastrophic forgetting, and enables balanced incremental learning by bridging the semantic gap between base and novel classes in the classifier space.
Figure 7: Examples of CUB-200 images under different noise conditions: (a) Original clean image, (b) Gaussian noise, (c) Snow, (d) Pixelate.
TABLE XIII: Performance of BiAG on CUB-200 under different noise conditions. Final Acc. represents the last session accuracy, and Avg. Acc. is the average across all sessions.
Noise Type Final Acc. (%) Avg. Acc. (%)
Original 63.72 70.71
Gaussian noise 59.58 68.74
Snow 55.84 62.40
Pixelate 56.28 64.72
Average (Noisy) 57.23 65.29
TABLE XIV: Comparison of model parameters and average accuracy on the CUB-200 dataset.
Method Params (M) Average Acc.
CEC (CVPR-2021) 12.34 61.33
LIMIT (ECCV-2022) 12.33 66.67
FACT (CVPR-2022) 11.34 64.42
NC-FSCIL (ICLR-2023) 15.91 67.28
ALICE (ECCV-2022) 41.71 65.75
SVAC (CVPR-2023) 24.29 69.35
YourSelf (ECCV-2024) 11.90 69.85
BiAG (Ours) 18.27 70.71
IV-G2 Analogical Reasoning
To intuitively demonstrate the analogical reasoning capability of our model, we visualize the semantic associations established by our generated weights. Specifically, we select two base classes—White Necked Raven and White Pelican—which exhibit distinct visual appearances and semantic meanings. For each class, we compute the similarity between its generated weight and all existing base class weights, and visualize the top-3 most similar classes.
As shown in Figure 6, the most similar classes for White Necked Raven include Groove Billed Ani, Brewer Blackbird, and Red Winged Blackbird, all of which share dark plumage and similar morphology. In contrast, the generated weight for White Pelican finds the closest matches in Laysan Albatross and California Gull, which share similar aquatic or seabird characteristics.
These examples indicate that our model is capable of capturing semantic-level analogies between novel and base classes, even when visual appearances vary significantly. This supports our claim that the generated weights are not arbitrary but stem from a reasoning process grounded in learned concepts.
TABLE XV: Inference efficiency on CUB-200.
Method Backbone Parameters (M) FLOPs (G) Final Acc.
Baseline-R18 ResNet18 12.34 1.82 61.81
BiAG (Ours) ResNet18 18.27 2.14 63.72
SVAC ResNet18 24.29 4.07 62.40
Baseline-R50 ResNet50 26.16 4.13 62.56
Baseline-R101 ResNet101 45.06 7.82 62.52
IV-H Robustness under Noisy Conditions
To further evaluate the robustness of BiAG, we conducted experiments on CUB-200 under multiple noisy conditions, as illustrated in Figure 7 and summarized in Table XIII. Noisy inputs represent a critical challenge in FSCIL settings because the limited number of novel-class samples makes the model particularly sensitive to input perturbations. BiAG achieves 70.71% accuracy on clean data and maintains an average of 65.29% under three different noise types, demonstrating that the proposed analogical weight generation framework preserves strong robustness even when inputs are corrupted.
IV-I Complexity Analysis
IV-I1 Parameter Comparison
To evaluate the efficiency of the proposed Brain-Inspired Analogical Generator (BiAG), we analyze the model complexity in terms of the total number of parameters and its relationship to average accuracy on the CUB-200 dataset. For a fair comparison, all methods are implemented using the ResNet-18 backbone.
Table XIV presents the comparison of model parameters and the corresponding average accuracy for several representative FSCIL methods. BiAG has 18.27M parameters, which is a moderate level of complexity compared with other approaches. Specifically, methods such as FACT, CEC, and LIMIT have 11.34M, 12.34M, and 12.33M parameters, respectively. While these models are lighter, they achieve lower average accuracy than BiAG. Conversely, SVAC and ALICE have 24.29M and 41.71M parameters, significantly larger than BiAG, yet offer comparable or lower accuracy. We also include the recent YourSelf baseline, which uses 11.90M parameters and attains 69.85% average accuracy on CUB-200. Compared to YourSelf, BiAG achieves +0.86% higher average accuracy. Our method demonstrates that moderate complexity is sufficient to achieve performance, balancing representational power and efficiency.
This analysis highlights that BiAG maintains high accuracy without incurring excessive parameter overhead, confirming its practicality for few-shot class-incremental learning scenarios. Future work will explore lightweight extensions of BiAG to further reduce complexity while preserving performance.
IV-I2 Training Efficiency
We evaluate the runtime and memory efficiency of BiAG on the CUB-200 dataset, as summarized in Table XII. BiAG completes the entire training process in 36.6 minutes, which is shorter than SVAC at 49.95 minutes and NC-FSCIL at 60.57 minutes, while requiring only 1.33 GB of memory in the incremental sessions. For YourSelf, since the authors release trained Base-session weights, its Base Training time is not applicable and is marked as “–” in Table XII. During incremental sessions, YourSelf requires 212.5 minutes and 9.51 GB of GPU memory, whereas BiAG needs only 2.10 minutes and 1.33 GB, respectively. Consequently, BiAG’s total time is 82.8% shorter than YourSelf.
Although an additional BiAG training stage is introduced, its incremental sessions require only 2.10 minutes each, enabling both efficient training and superior accuracy of 70.71%. These results indicate that BiAG achieves a favorable balance between computational cost and recognition performance, supporting its practicality for FSCIL tasks. All methods are profiled on a single NVIDIA RTX 3090 GPU to ensure fair timing and memory measurements.
IV-I3 Inference Efficiency and Analogical Learning
To evaluate whether BiAG’s performance improvements are primarily attributed to its analogical learning mechanism rather than the additional parameters introduced by WSA and WPAA, we analyze parameter efficiency and inference performance across different backbone configurations. The results on the CUB-200 dataset are summarized in Table XV.
Our BiAG with ResNet-18 achieves 63.72% final accuracy with only 18.27M parameters and 2.14G FLOPs, outperforming larger backbones such as ResNet-50 (62.56%) and ResNet-101 (62.52%), as well as the SOTA method SVAC (62.40%). This confirms that the performance gain comes primarily from the proposed analogical weight generation mechanism, which effectively leverages semantic relationships between base and novel classes, rather than relying on simply increasing model size or inference cost.
@@ V Conclusion and Future Work
We proposed the Brain-Inspired Analogical Generator (BiAG), a novel framework for Few-Shot Class-Incremental Learning (FSCIL). Inspired by human analogical reasoning, BiAG generates novel class weights by establishing semantic correspondences with learned knowledge, avoiding parameter updates in incremental sessions. Leveraging old class weights, new prototypes, and a learnable query through its WPAA, WSA, and SCM modules, BiAG achieves strong generalization and reduced forgetting. Extensive experiments on CIFAR-100, miniImageNet, and CUB-200 demonstrate performance across all key metrics.
Future work will explore extending analogical generation to domain-incremental and open-vocabulary settings, as well as integrating BiAG with advanced backbones or vision-language models like CLIP. We believe this brain-inspired approach offers a promising direction for flexible and cognitively motivated continual learning systems.